OpenStudy (anonymous):

If y varies directly as x, what is the value of y in these ordered pairs? (3, 12) and (24, y)

OpenStudy (callisto):

\[y \ \alpha \ x\]\[\frac{y_1}{x_1} =\frac{y_2}{x_2}\]\[\frac{12}{3} =\frac{y}{24}\]Can you solve y?

OpenStudy (anonymous):

y=6?

OpenStudy (callisto):

Please read the question: y varies directly as x

OpenStudy (lgbasallote):

exactly why i said sorry

OpenStudy (lgbasallote):

i'll wear my glasses now...

OpenStudy (callisto):

@emayorga y is not 6 .... y = 24 x 12 / 3 = ?